OpenStudy (anonymous):

If a Canada goose completes a 2400-mile migration in exactly 13 days and 8 hours, then what was her average speed for the whole trip in miles per hour?

OpenStudy (anonymous):

Does any one have an answer

OpenStudy (anonymous):

miles per hour miles / hour \[\frac{2400}{13\ 24+8}=\frac{15}{2}=7.5 \text{ mph} \]13 24 is 13 times 24

OpenStudy (anonymous):

30 mph

OpenStudy (anonymous):

13 days and 8 hours is 80 hours. So you would do 2400 miles traveled divided by 80 hours to get 30 miles per hour

OpenStudy (anonymous):

There is 24 hours in one day. 13 days is 13*24 or 312 hours.

OpenStudy (anonymous):

oh I'm sorry